# A Hamiltonian for 1:1 Rotational Secondary Resonances, and Application to Small Satellites of Saturn and Jupiter

### v1

@@ 2.2 Perturbation of the integrable hamiltonian
where $H^{\prime}_{0}$ is given in (7) and $H^{\prime}_{1}$ is the perturbation. We will include in $H^{\prime}_{1}$ the two neighboring resonances of the synchronous resonances, namely, 3:2 and 1:2 (see Equation (4)). In terms of old variables $(\phi,\Phi)$ we can obtain from (4) that:
@@
After substituting the expression (11) in (17) and the result in (16), we can obtain the perturbation $H^{\prime\prime}_{1}(\psi,J,t)$ as follows:
@@
$+$ $(3/16)F_{1}^{2}\cos{nt}+(3/16)F_{3}^{2}\cos{nt}-(3/4)\cos{nt}$
-- $(1/2)F_{1}\cos{(nt-\psi)}$
$+$ $(1/16)F_{1}^{3}\cos{(nt-\psi)}$
-- $(1/16)F_{1}^{2}F_{3}\cos{(nt-\psi)}$
$+$ $(1/8)F_{1}F_{3}^{2}\cos{(nt-\psi)}$
-- $(3/32)F_{1}^{2}\cos{(nt-2\psi)}$
$+$ $(3/16)F_{1}F_{3}\cos{(nt-2\psi)}$
$+$ $(1/16)F_{3}^{3}\cos{(nt-3\psi)}$
-- $(1/48)F_{1}^{3}\cos{(nt-3\psi)}$
$+$ $(1/8)F_{1}^{2}F_{3}\cos{(nt-3\psi)}$
-- $(1/2)F_{3}\cos{(nt-3\psi)}$
-- $(3/16)F_{1}F_{3}\cos{(nt-4\psi)}$
$+$ $(1/16)F_{1}F_{3}^{2}\cos{(nt-5\psi)}$
-- $(1/16)F_{1}^{2}F_{3}\cos{(nt-5\psi)}$
-- $(3/32)F_{3}^{2}\cos{(nt-6\psi)}$
-- $(1/16)F_{1}F_{3}^{2}\cos{(nt-7\psi)}$
-- $(1/48)F_{3}^{3}\cos{(nt-9\psi)}$
-- $(1/16)F_{1}^{3}\cos{(nt+\psi)}$
$+$ $(1/2)F_{1}\cos{(nt+\psi)}$
-- $(1/8)F_{1}F_{3}^{2}\cos{(nt+\psi)}$
$+$ $(1/16)F_{1}^{2}F_{3}\cos{(nt+\psi)}$
$+$ $(3/16)F_{1}F_{3}\cos{(nt+2\psi)}$
-- $(3/32)F_{1}^{2}\cos{(nt+2\psi)}$
-- $(1/16)F_{3}^{3}\cos{(nt+3\psi)}$
$+$ $(1/48)F_{1}^{3}\cos{(nt+3\psi)}$
$+$ $(1/2)F_{3}\cos{(nt+3\psi)}$
-- $(1/8)F_{1}^{2}F_{3}\cos{(nt+3\psi)}$
-- $(3/16)F_{1}F_{3}\cos{(nt+4\psi)}$
$+$ $(1/16)F_{1}^{2}F_{3}\cos{(nt+5\psi)}$
-- $(1/16)F_{1}F_{3}^{2}\cos{(nt+5\psi)}$
-- $(3/32)F_{3}^{2}\cos{(nt+6\psi)}$
$+$ $(1/16)F_{1}F_{3}^{2}\cos{(nt+7\psi)}$
$+$ $(1/48)F_{3}^{3}\cos{(nt+9\psi)}.$
Note that at this step we are not utilizing the expanded versions of $F_{1}$ and $F_{3}$ given in the right side of Equations (12) and (13).
@@ 3.1 The 1:1 secondary resonance
we can obtain the disturbing part of the hamiltonian associated with the 3:1 secondary resonance utilized in the theory developed by Wisdom (2004) to study the rotation of Enceladus (see Section 1).
The problem we are considering in this work is related to the nature of the $\alpha$ and $\beta$ regimes of motion in the phase space of rotation of Methone and, as pointed out in Section 1, we conjectured that the coexistence of the regimes is explained by a 1:1 secondary resonance. We will prove our conjecture in the next two subsections after considering the terms factored by $\cos{(\psi-nt)}$ in (18) so that we have the following disturbing part of the hamiltonian:
$H^{\prime\prime}_{1}(\psi,J,t)=-\frac{\epsilon^{2}n^{2}Ce}{4}\left(-\frac{1}{4}F_{1}^{2}F_{3}-2F_{1}+\frac{1}{4}F_{1}^{3}+\frac{1}{2}F_{1}F_{3}^{3}\right)\cos{(\psi-nt)}.$ (20)
@@ 3.2 The second extension of the phase space and the non-singular variables
$\psi^{\prime}$ $=$ $\psi-nt,$ (21)
$J^{\prime}$ $=$ $J,$
and define
@@
where $\hat{J}$ is a dimensionless quantity (Wisdom 2004), and $\delta$ can be interpreted as being the distance of the exact resonance $\frac{1}{1}$ secondary resonance.
After collecting the terms in (20) up to order $J^{\prime 2}$, neglecting the constant terms, and factoring the parameter $n^{2}C$, we have the final form in Angle-Action variables of the hamiltonian (12) for 1:1 secondary resonance:
$H^{\prime\prime\prime}_{1:1}(\psi^{\prime},\hat{J},t)=n^{2}C\left[\delta\hat{J}-\frac{\hat{J}^{2}}{4}+\frac{3}{8}e\sqrt{\epsilon}\left(\sqrt{2\hat{J}}\right)^{3}-e\sqrt{\epsilon^{3}}\sqrt{2\hat{J}}\right]\cos{\psi^{\prime}}.$ (23)
@@
$H\equiv\frac{H^{\prime\prime\prime}_{1:1}}{n^{2}C}.$ (24)
@@ 3.3 Atlas of the phase space
Keeping in hand a time-independent one-degree-of-freedom hamiltonian for our problem, we can now explore the rotation phase space by computing the level curves of Equation (26). We also aim to compare the level curves with the surfaces of sections and a suitable plane of variables which can be the same utilized in Fig. (1), $(\theta,\frac{\dot{\theta}}{n})$. To relate them to the variables $(x,y)$ (Equation (25)), we first recall that in the linear approximation of harmonic oscillator the relation between variables $(\phi,\Phi)$ (Equation (6)) and the Angle-Action $(\psi,J)$ are given by the canonical transformation
$\phi$ $=$ $\left(\frac{2J}{\alpha}\right)^{1/2}\sin{\psi},$
$\Phi$ $=$ $(2J\alpha)^{1/2}\cos{\psi},$ (27)
$\alpha$ $=$ $\epsilon nC.$
By fixing a instant of time ($t=0$ for instance), manipulating the Equations (6), (25), (27), and by noting that at $t=0$, $\psi^{\prime}=\psi$ (Equation (21)), we can show that,
@@
Fig. 2 shows the level curves of the hamiltonian (26) for six values of $\epsilon$. Let $\epsilon_{c}$ be a critical value of $\epsilon$. For $\epsilon<\epsilon_{c}$, only the equilibrium point associated with the $\alpha$-regime exists. The rising of the $\beta$-regime occurs at values of $\epsilon$ slightly larger than $\epsilon_{c}$, and now three equilibria centers share the $y$-axis: two of them are stable points, namely those associated to $\alpha$ and $\beta$-regimes, and the third one is an unstable equilibrium point associated to the separatrix of the $\beta$-regime. The plot in bottom-middle panel in Fig. 2 reproduces with good agreement the surface of section given in Fig. 1. Next we will study in detail the bifurcation of the regimes in the phase space and determine $\epsilon_{c}$ numerically.
Figure 2: Contour plots of the hamiltonian (26) for several values of the non-sphericity parameter $\epsilon$. The orbital eccentricity is $0.0014$ and the range of values of $\epsilon$ are compatible with the values of Methone ($\epsilon\sim 1.077$ (the same utilized in Fig. 1). Green and blue lines correspond to level curves within the domains of the $\alpha$ and $\beta$-regimes, respectively. Red curves are the separatrix of the $\beta$-regime. Right-bottom plot is a detailed close of the fixed point associated with the $\beta$-regime. The dashed line indicates the forced component.
3.4 The forcing of the $\beta$-regime and the quasi-synchronous regime.
The loci of the equilibria centers (fixed points) of the $\alpha$ and $\beta$-regimes in the phase space can be obtained analytically from equations of motion:
@@
The equilibria conditions $(\frac{dx}{dt},\frac{dy}{dt})=(0,0)$ lead to a system of non-linear algebraic equations, and equilibria solutions in the x-axis where $y=0$ must satisfy the equation:
@@
The three solutions of Equation (31) are only real for values of $\epsilon$ larger than $\epsilon_{c}$. For $\epsilon<\epsilon_{c}$, only one root is real and there is only one fixed point in the phase space, namely, that one related to the $\alpha$-regime. At $\epsilon=\epsilon_{c}$ occurs the rising of the $\beta$-regime and the unstable point. Considering the current $\epsilon=1.077$ of Methone, the solution of Equation (31) is $(x_{1},x_{2},x_{3})=(-0.5616685,0.02034028,0.5478663)$. Substituting these roots into Equation (28) we obtain $\frac{\dot{\theta}}{n}=0.4171$ ($x_{1}$: $\alpha$-regime); $\frac{\dot{\theta}}{n}=1.0211$ ($x_{2}$: $\beta$-regime); and $\frac{\dot{\theta}}{n}=1.5686$ ($x_{3}$: separatrix). The three values agree with those obtained after inspection of the level curves given Fig. 2. Fig. 3 shows the real solutions calculated in this way in the range of $1.080.96<\epsilon<$ 1.08. We can see that the bifurcation occurs at $\epsilon_{c}\sim 1.0149$. Note also the quasi-symmetry of the roots $x_{1}$ and $x_{3}$.
The fixed point of the $\beta$-regime is not centered at the origin as it is indicated by the dashed line in Fig. (2), bottom-right. The forced component calculated above deviates from the unit such that $\frac{\dot{\theta}}{n}=1.0211$. More generally note that in the case of non-circular orbits, $x=0$ is never a solution of Equation (31), so that from Equation (28) the forcing is never null in this case (recall the definition of $\gamma_{1}$).
@@
Figure 3: The real roots of Equation (31) as function of $\epsilon$. Red, blue and green curves are the loci of the fixed points of the separatrix ($s$), the $\beta$, and the $\alpha$-regimes, respectively. Dashed horizontal line, dashed-dot-vertical line, and dashed-dot-dot vertical line indicate the straight $\frac{\dot{\theta}}{n}=1$, the critical value $\epsilon_{c}=1.0149$ where the $\beta$-regime rises, and the current value for Methone ($\epsilon=1.077$), respectively.

### v3

@@ 2.2 Perturbation of the integrable hamiltonian
where $H^{\prime}_{0}$ is given in (7) and $H^{\prime}_{1}$ is the perturbation. We will include in $H^{\prime}_{1}$ the two neighboring resonances of the synchronous resonances, namely, 3:2 and 1:2, which are proportional to $e$ (see Equation (4)). In terms of old variables $(\phi,\Phi)$ we can obtain from (4):
@@
After substituting the expression (9) in (15) and the result in (14), we can obtain the perturbation $H^{\prime\prime}_{1}(\psi,J,t)$ as follows [^5] The addition of the next term in the expansion of Equation (9) does not generate additional terms of order of $J^{3}$ in the cases of the resonances 3/1 and 1/1 considered in this work. The same is not necessarily true in other cases, like the 5/1 and 6/1 secondary resonances.:
@@
$[(3/16)F_{1}^{2}+(3/16)F_{3}^{2}-3/4]\cos{nt}+$
$[(-1/2)F_{1}+(1/16)F_{1}^{3}-(1/16)F_{1}^{2}F_{3}+(1/8)F_{1}F_{3}^{2}]\cos{(nt-\psi)}+$
$[(-3/32)F_{1}^{2}+(3/16)F_{1}F_{3}]\cos{(nt-2\psi)}+$
$[(1/16)F_{3}^{3}-(1/48)F_{1}^{3}+(1/8)F_{1}^{2}F_{3}-(1/2)F_{3}]\cos{(nt-3\psi)}+$
$((-3/16)F_{1}F_{3}\cos{(nt-4\psi)}+$
$[(1/16)(F_{1}F_{3}^{2}-F_{1}^{2}F_{3})]\cos{(nt-5\psi)}+$
$(-3/32)F_{3}^{2}\cos{(nt-6\psi)}+$
$(-1/16)F_{1}F_{3}^{2}\cos{(nt-7\psi)}+$
$(-1/48)F_{3}^{3}\cos{(nt-9\psi)}+$
$[(-1/16)F_{1}^{3}+(1/2)F_{1}-(1/8)F_{1}F_{3}^{2}+(1/16)F_{1}^{2}F_{3}]\cos{(nt+\psi)}+$
$[(3/16)F_{1}F_{3}-(3/32)F_{1}^{2}]\cos{(nt+2\psi)}+$
$[(-1/16)F_{3}^{3}+(1/48)F_{1}^{3}+(1/2)F_{3}-(1/8)F_{1}^{2}F_{3}]\cos{(nt+3\psi)}+$
$(-3/16)F_{1}F_{3}\cos{(nt+4\psi)}+$
$[(1/16)(F_{1}^{2}F_{3}-F_{1}F_{3}^{2})]\cos{(nt+5\psi)}+$
$(-3/32)F_{3}^{2}\cos{(nt+6\psi)}+$
$(1/16)F_{1}F_{3}^{2}\cos{(nt+7\psi)}+$
$(1/48)F_{3}^{3}\cos{(nt+9\psi)}.$
@@ 3.1 The 1/1 secondary resonance
we can obtain the disturbing part of the hamiltonian associated with the 3/1 secondary resonance utilized in the theory developed by Wisdom (2004) to study the rotation of Enceladus (see Section 1) [^6] The last two terms in Equation (17) are higher order and don’t contribute to Wisdom’s model of the 3/1 secondary resonance, which is valid up to $J^{2}$..
The problem we are considering in this work is related to the nature of the $\alpha$ and $\beta$-regimes of motion in the phase space of rotation of Methone and other satellites; as pointed out in Section 1, we conjectured that the coexistence of the regimes is explained by a 1/1 secondary resonance. We will prove our conjecture in the next two subsections after considering the terms factored by $\cos{(\psi-nt)}$ in (16) so that we have the following disturbing part of the hamiltonian:
$H^{\prime\prime}_{1}(\psi,J,t)=-\frac{\epsilon^{2}n^{2}Ce}{4}\left(+\frac{1}{4}F_{1}^{2}F_{3}+2F_{1}+\frac{1}{4}F_{1}^{3}-\frac{1}{2}F_{1}F_{3}^{2}\right)\cos{(\psi-nt)}.$ (18)
From Equations (10) and (11), keeping the terms up to $O(J^{2})$, after neglecting the constant terms we obtain:
$H^{\prime\prime}_{1}(\psi,J,t)=e\left[+\frac{3}{8}\sqrt{\frac{\epsilon n}{C}}\left(\sqrt{2J}\right)^{3}-\sqrt{\frac{\epsilon^{3}n^{3}}{C}}\sqrt{2J}\right]\cos{(\psi-nt)}.$ (19)
@@ 3.2 The second extension of the phase space and the non-singular variables
Including terms up to $J^{2}$ in Equation (8) and discarding constants terms we have
$H^{{}^{\prime\prime}}_{0}(-,J)=\epsilon nJ-\frac{1}{4C}J^{2}.$ (20)
@@
$\psi^{\prime}$ $=$ $\psi-nt,$
$J^{\prime}$ $=$ $J.$ (21)
From (20) and (19) the full hamiltonian $H^{\prime\prime}_{1/1}(\psi^{\prime},J^{\prime},t)=H^{{}^{\prime\prime}}_{0}(J^{\prime})+H^{\prime\prime}_{1}(\psi^{\prime},J^{\prime},t)$ becomes
$H^{\prime\prime\prime}_{1/1}(\psi^{\prime},\hat{J})=-nJ^{\prime}+\epsilon nJ^{\prime}-\frac{1}{4C}J^{\prime 2}+e\left[+\frac{3}{8}\sqrt{\frac{\epsilon n}{C}}\left(\sqrt{2J^{\prime}}\right)^{3}-\sqrt{\frac{\epsilon^{3}n^{3}}{C}}\sqrt{2J^{\prime}}\right]\cos{\psi^{\prime}},$
where the term $-nJ^{\prime}$ comes from a phase space extension.
Define
@@
where $\hat{J}$ is a dimensionless quantity (Wisdom 2004), and $\delta$ can be interpreted as being the distance of the exact resonance $/1$ 1 secondary resonance.
Factoring the parameter $n^{2}C$, we can obtain after trivial algebra the full hamiltonian in action-angle variables:
$H^{\prime\prime\prime}_{1/1}(\psi^{\prime},\hat{J})=n^{2}C\left[\delta\hat{J}-\frac{\hat{J}^{2}}{4}+e\left(\frac{3}{8}\sqrt{\epsilon}\left(\sqrt{2\hat{J}}\right)^{3}-\sqrt{\epsilon^{3}}\sqrt{2\hat{J}}\right)\cos{\psi^{\prime}}\right].$ (23)
@@
$H\equiv\frac{H^{\prime\prime\prime}_{1/1}}{n^{2}C}.$ (24)
@@ 3.3 Atlas of the phase space
Keeping in hand a time-independent one-degree-of-freedom hamiltonian for our problem, we can now explore the rotation phase space by computing the level curves of Equation (26). We also aim to compare the level curves with the surfaces of sections and a suitable plane of variables which can be the same utilized in Fig. 1, $(\theta,\frac{\dot{\theta}}{n})$. To relate them to the variables $(x,y)$ (Equation (25)), we first recall that in the linear approximation of harmonic oscillator the relation between variables $(\phi,\Phi)$ (Equation (6)) and the angle-action $(\psi,J)$ are given by the canonical transformation
$\phi$ $=$ $\left(\frac{2J}{\lambda}\right)^{1/2}\sin{\psi},$
$\Phi$ $=$ $(2J\lambda)^{1/2}\cos{\psi},$ (27)
where $\lambda=\epsilon nC$. By fixing a instant of time ($t=0$ for instance), manipulating the Equations (6), (25), (27), and by noting that at $t=0$, $\psi^{\prime}=\psi$ (Equation (21)), we can show that,
@@
Fig. 2 shows the level curves of the hamiltonian (26) for six values of $\epsilon$. Let $\epsilon_{c}$ be a critical value of $\epsilon$. For $\epsilon<\epsilon_{c}$, only the equilibrium point associated with the $\alpha$-regime exists. The rising of the $\beta$-regime occurs at values of $\epsilon$ slightly larger than $\epsilon_{c}$, and now three equilibria centers share the $y$-axis: two of them are stable points, namely those associated to $\alpha$ and $\beta$-regimes, and the third one is an unstable equilibrium point associated to the separatrix of the $\beta$-regime. The plot in right-bottom panel in Fig. 2 reproduces with good agreement the surface of section given in Fig. 1. The two curves given in blue and purple superposed to the level curves in the plot $\epsilon=1.077$ are the surfaces of section. The agreement is good, in spite small of differences in the position of the fixed points. The model can be improved by including high-order terms in the expansions in action to obtain better agreements (see Section 4).
Some investigations on the dependence of eccentricity of Methone on the phase space is given in Appendix.
Next we will study in detail the bifurcation of the regimes in the phase space and determine $\epsilon_{c}$ numerically.
Figure 2: Contour plots of the hamiltonian (26) for several values of the non-sphericity parameter $\epsilon$. The orbital eccentricity is $0.0014$ and the range of values of $\epsilon$ are compatible with the values of Methone, $\epsilon\sim 1.077$ (the same utilized in Fig. 1). Green and blue lines correspond to level curves within the domains of the $\alpha$ and $\beta$-regimes, respectively. Red curves are the separatrix of the $\beta$-regime. Right-bottom plot is a detailed close of the fixed point associated with the $\beta$-regime. The dashed line indicates the forced component. The two curves given in blue and purple superposed to the level curves in the plot $\epsilon=1.077$ are the surfaces of section of the hamiltonian (1).
Figure 3: (a) The real roots of Equation (31) as function of $\epsilon$. Red, blue and green curves are the loci of the fixed points of the separatrix ($s$), the $\beta$ and the $\alpha$-regimes, respectively. Dashed horizontal, dashed-dot vertical, and dashed-dot-dot vertical lines indicate, respectively, the straight $\frac{\dot{\theta}}{n}=1$, the critical value $\epsilon_{c}=1.0149$ where the $\beta$-regime rises, and the current value for Methone ($\epsilon=1.077$). Horizontal lines give the values at the corresponding y-axes. (b) Amplitude of $\phi=\theta(t)-nt$ close to fixed point of the $\alpha$-regime. (c) The same as (a) in the case of $\beta$-regime. (d) The curve of $\frac{\dot{\theta}}{n}=x\sqrt{\epsilon}+1$ in the case of the root $x_{2}$.
3.4 The forcing of the quasi-synchronous $\beta$-regime
The loci of the equilibria centers (or the fixed points in the full problem) of the $\alpha$ and $\beta$-regimes in the phase space can be obtained analytically from equations of motion:
@@ 3.4 The forcing of the quasi-synchronous $\beta$-regime
The equilibrium condition $(\frac{dx}{dt},\frac{dy}{dt})=(0,0)$ leads to a system of non-linear algebraic equations, and equilibria solutions in the x-axis where $y=0$ must satisfy the equation:
@@
The three solutions of Equation (31) are real only for values of $\epsilon$ larger than $\epsilon_{c}$. For $\epsilon<\epsilon_{c}$, only one root is real and there is only one fixed point in the phase space, namely, that one related to the $\alpha$-regime. At $\epsilon=\epsilon_{c}$ occurs the rising of the $\beta$-regime and the unstable point. Considering the current $\epsilon=1.077$ of Methone, the solution of Equation (31) is $(x_{1},x_{2},x_{3})=(-0.5616685,0.02034028,0.5478663)$. Substituting these roots into Equation (28) we obtain $\frac{\dot{\theta}}{n}=0.4171$ ($x_{1}$: $\alpha$-regime); $\frac{\dot{\theta}}{n}=1.0211$ ($x_{2}$: $\beta$-regime); and $\frac{\dot{\theta}}{n}=1.5686$ ($x_{3}$: separatrix). The three values agree with those obtained after inspection of the level curves given Fig. 2. Fig. 3a) shows the real solutions calculated in this way in the range of $1.080.96<\epsilon<$ 1.08. We can see that the bifurcation occurs at $\epsilon_{c}\sim 1.0149$. Note also the quasi-symmetry of the roots $x_{1}$ and $x_{3}$.
The fixed point of the $\beta$-regime is not centered at the origin as it is indicated by the dashed line in Fig. 2, bottom-right. The forced component calculated above deviates from the unit such that $\frac{\dot{\theta}}{n}=1.0211$. More generally, note that in the case of non-circular orbits, $x=0$ is never a solution of Equation (31), so that from Equation (28) the forcing is never null in this case.
@@
3.5 Amplitude of the physical libration
We can obtain an approximate expression of the amplitude of libration of the angle $\phi=\theta(t)-nt$ from equation (24) in Wisdom (2004):
$S=\frac{1}{2}\left(8\frac{\hat{J}}{\epsilon}+4\frac{\hat{J}^{2}}{\epsilon^{2}}\right)^{1/2}.$ (33)
The next step is to calculate the amplitude of libration of $\phi$ as a function of $\epsilon$. For this task, we need to evaluate the action $\hat{J}$ in the equation above, and following Wisdom’s method, we utilize the stable equilibria points of the hamiltonian. At equilibrium $x=\sqrt{2\hat{J}}$, so that the action of the stable equilibrium is $\hat{J}=\frac{x^{2}}{2}$, where $x$ is the corresponding solution of Equation (31).
The plot of $S$ in the case of the secondary resonance (denoted by $S_{\alpha}$) as function of $\epsilon$ is given in Fig. 3b). The amplitude of $\phi=\theta(t)-nt$ increases to values close to the current one indicated by vertical and horizontal lines. There is good agreement when compared to the fixed point in Fig. 1c) and the equilibria points given in Fig. 2.
In Fig. 3(c) it is shown the numerical value of $S$ is the case of the $\beta$-regime (denoted by $S_{\beta}$). In the case the amplitude of $\phi=\theta(t)-nt$ it is always smaller than that case of the $\alpha$-regime. Note that $S_{\beta}$ decreases as $\epsilon$ increases. This is related to the fact the forced component is larger for values of $\epsilon$ closer to the $\epsilon_{c}$ as seen in Fig. 3d), where $\frac{\dot{\theta}}{n}=x\sqrt{\epsilon}+1$ is plotted in the case of the root $x_{2}$. Note the good agreement with the fixed points given in Fig. 2.
The amplitude of variation of the $\phi=\theta(t)-nt$ in the $\alpha$-regime is much larger than in the case of the $\beta$-regime. The current rotation state of the regular satellites of the Solar System is the synchronous equilibrium (e.g. Peale 1977). It is well known that this final stage of rotation has been reached after a long-term evolution due to planet-satellite tidal interactions (see Wisdom 2004, Ferraz-Mello 2013, and references therein). The corresponding amplitudes of the physical libration are always very small; Enceladus has one of the largest values $\sim 0.12$ degree. Therefore, the $\beta$-regime is the probable final state of rotation of these sorts of bodies.
In the case of small satellites, the existence of secondary resonances due to their irregular shapes leads to the question of the possibility of high-amplitude libration on $\alpha$-regime. For instance, in the case of Saturnian small satellites, the most recent observations and analyses give small amplitudes of libration (Thomas and Helfenstein 2020, Rambaux et al. 2022). According to Melnikov and Schevchenko (2022), the amplitude of libration of Amalthea (Jupiter) is $<5$ degree and would be $\sim 30$ degree if the $\alpha$-regime had been attained (Pashkevich et al. 2021). Therefore, based on observations, it is probable that, in despite of their large $\epsilon$, all small satellites of Saturn and Jupiter also currently rotate in the equilibrium site associated with the $\beta$-regime.